Determine which two equations represent parallel lines. Show work.
(a) y=6x-5 (b) y=-6x+5 (c) y=6x+1 (d) y=-1/6x-5

Found 2 solutions by Nate, sister_insomnia:
Answer by Nate(3500)

They are all in slope-intercept form: y = mx + b
m is the slope .... so find equations that m is the exact same
(a) and (c)

Answer by sister_insomnia(8)

(a) y=6x-5 (b) y=-6x+5 (c) y=6x+1 (d) y=-1/6x-5
To determine which two represent parallel lines is easy because your equations are in slope-intercept form. y=mx+b in which m is the slope:
So, (a) Has a slope of 6, (b) Has a slope of -6, (c) Has a slope of 6, and (d) Has a slope of -1/6. When concerning yourself with parallel lines remember that the lines have the same slope. Therefore, a and c, which both have a slope of 6, are your answers.

Good Luck!
S_I